The Student Information Processing Board (SIPB)

• Who are we?

The Student Information Processing Board (SIPB, pronounced sip-bee) is MIT’s volunteer student group dedicated to computing. We have an office in the MIT Student Center, and you can come ask us for help if you get stuck on anything computer-related.

• What do we do?

We make it easier for you to use computers at MIT. That may sound kind of boring, but it’s actually a lot of fun, as well as being tremendously important. SIPB picks up the slack when the official MIT computing infrastructure is failing to do something we think is important. Most of the time, we just hang out in the office, doing our own thing and helping people who stop by or call us with questions. We also develop and improve software (some of our members are integral to the Linux, NetBSD and GNOME development communities,) make it easier for students to access Athena from their dorms, and teach classes over IAP about programming and maintaining computers. We’re always looking for more people to help with these things!

• Why should you join? The best reason to join SIPB is that you are interested in learning about and helping people with computers. If you’ve been using computers for a long time and enjoy talking about them, you’re going to love SIPB. But, all that’s really necessary to join SIPB is an interest in and curiosity about computers. Amongst our members are some of the most competent and influential programmers in the world, and while we try to help any MIT student, we help each other with technical issues even more. You’d be amazed at how much you can learn just sitting in the office and listening to the conversations around you. The SIPB office is also a great place to hack. It’s a lot more fun than hacking alone in your dorm room. And no one here will ever tell you that your conversation is too geeky.

• How do you join?

We have a two-tier membership system due to the fact that we have privileged access to a huge array of hardware and some socially important MIT network services. Anyone can become a “prospective member” meaning you have access to SIPB’s resources. After you have been around a while, and there’s a perception that what you’re doing is useful to SIPB, then you can become a “member”, which means you get your own key to the office and full administrative access to most of our machines. Official meetings are Monday at 7:30 p.m. in the office. You must attend four meetings before you can become a member. Before and after a meeting there are usually more members and prospectives around than at other times, but you can hang around the office any time it’s open, and it’s often more fun at other times.

• How can you reach us? The SIPB office is W20-557, on the fifth floor of the Student Center. Feel free to drop by and ask about joining. The office doesn’t have set hours, but it’s almost always open between 1pm and 3am. You can also reach us by calling x3-7788 or sending mail to sipb@mit.edu.

SIPB Project List

A lot of what SIPB does which is not helping users with immediate problems and answering questions falls under the heading of SIPB projects. SIPB projects are usually things that a member or two have decided are worth doing and are working on. The main qualification of SIPB projects is that they help the MIT community (or maybe just the SIPB community) in some way, generally involving computing.

If you have any ideas for projects that you’d like to do, or would like help with, or just think would be a good idea for somebody to do, we’d love to hear about it. SIPB is a great resource for both tools and knowledge. If you don’t have your own idea, below is a list of ongoing and proposed projects, with their contact address that you are welcome to join in on.

Webmaster - webmaster In charge of running the web server at www.mit.edu and the services that run on it. People on webmaster also answer questions that are sent in by people using various websites.

Locker Software - sipb SIPB maintains a bunch of software stored in software “lockers”. New people interested in maintaining pet software or in upgrading aging software are always welcome.

Usenet - usenet SIPB runs the primary MIT Usenet news server (news.mit.edu). Documents - aletta A lot of things around MIT and SIPB are just poorly documented, and

would be really useful to people if they were documented. SIPB maintains some documents - they are kept in bins next to the door to the office - and there are certainly others which could stand to be written. Some of the existing ones are online at http://www.mit.edu/sipb/docs.html.

• The Inessential Guide to Athena is currently being updated. We need people to write all different sections. This is a good way for you to learn about Athena, and something desperately needed by the MIT community.

• SIPB Office Manual is another document that needs updating. It contains information about SIPB’s projects, how it operates, and how to become a member. It also includes a brief history of SIPB. Contact lgdean@mit.edu if you’re interested.

• Accessing AFS from Windows

Classes - sipb-iap-sucker Many SIPB members teach classes during IAP and other times. A list

of the most recent IAP classes is available at http://websis.mit.edu/iap/nssipb.html. Debian Installer - sipb-debian The SIPB Debian project is working to create an installer that

combines the flexibility of Debian GNU/Linux with access to Athena features including MIT mail, sharing files via AFS and communication via Zephyr. Currently we have working code with some known problems.

Linux Installer - linux-dev An installer for Linux-Athena that works on non-cluster hardware. Based on the RedHat linux installer. Good for personal and lab workstations.

Other projects that might be cool for someone to do. Email sipb@mit.edu if you want to work on one of them:

• Install some decent ICQ clients, make the im locker have clients for other major IM systems.

• NetBSD Sparc.

• DUX Athena. Athena software running on Digital Unix, specifically on Alpha workstations. There is one machine in the SIPB office running this port of Athena, but it has not escaped yet. - ducks@mit.edu

• MacOS X. Is it Unix? What can we do with it? Currently we have a machine running OS X in the office. It has AFS and Kerberos and developer tools, but doesn’t really have any direction in its life.

• Give the Oracle server the love it needs. We have an Oracle server and Oracle wizards who would be happy to help.

• Allow user to build active pages (i.e. CGI or java servlets or something) on www or some other server. The security problem is an interesting one, as is making everything automatic.

• Web message forum for use by classes/others.